# Mapping the Trust Terrain: LLMs in Software Engineering - Insights and Perspectives

### v1

@@ 7.4. Discussion
Our study highlights the need for a comprehensive framework to measure trust in LLM4SE, as existing approaches fail to account for its multifaceted nature. While technical metrics such as accuracy, robustness, and accountability are crucial, our survey results indicate that trust also depends on usability, comprehensibility, and ethical considerations. Developers emphasized that trust is not solely determined by correctness but also by how seamlessly LLMs integrate into their workflows. The need for minimal debugging and rework reinforces the importance of trust metrics that capture practical usability alongside model performance.
Survey findings also emphasize that trust measurement should be multidimensional. A strong preference (96% of participants) for multi-item trust metrics aligns with broader research indicating that single-item trust measures oversimplify complex trust dynamics. This is particularly critical in SE, where small errors can lead to security vulnerabilities and software failures. A multidimensional approach—incorporating accuracy, usability, ethical compliance, and consistency—provides a more holistic assessment of trustworthiness.
Another major gap in current trust measurement frameworks is their failure to capture the temporal evolution of trust. Our findings indicate that trust in LLMs fluctuates over time, shaped by output quality and the extent to which developers feel they can rely on generated suggestions. This aligns with research in robotics and education, where trust is measured longitudinally through failure rates and behavioral consistency  (Lee and See, 2004b; Lewis et al., 2018). Future SE trust assessment frameworks should move beyond static evaluations and incorporate temporal metrics that track trust development over prolonged interactions with LLMs.
@@ 7.5. Opportunities for Future Work
The results of this study present several avenues for future research aimed at enhancing the metrics for trust in LLM4SE. Trust in LLMs is influenced by a variety of factors, including model accuracy, usability, ethical compliance, interpretibility, and robustness. Future research should focus on developing multi-item trust measurement frameworks that assess these dimensions in the context of SE. For instance, accuracy could be evaluated through correctness, alignment with user intentions, and contextual appropriateness. Similarly, usability could include measures such as integration ease, user control, and the effort required for debugging. Ethical compliance should be assessed through fairness evaluations and alignment with professional coding standards. A robust, multi-dimensional framework would provide a comprehensive understanding of trust and guide model improvements.
@@
Trust in LLMs evolves over time, influenced by repeated interactions and exposure to the model. Future research should explore longitudinal studies that track how trust builds or erodes as users gain experience with the model. These studies could examine changes in trust based on output quality, model performance over time, and user feedback. Understanding the temporal dynamics of trust would help in designing models that foster long-term user confidence, making them more effective and reliable for continuous use in SE tasks.
@@
Context-specific trust metrics are another crucial area for future work. Trust may be prioritized differently in tasks such as code generation, program repair, or test case generation, where each task involves distinct expectations and challenges. For example, code generation may emphasize accuracy, ethical considerations, and contextual relevance, while program repair may require metrics that account for error minimization, maintainability, and adherence to coding standards. Tailored trust metrics will ensure that the measurement framework aligns with the specific needs of each task and provides more meaningful insights into trust.
@@
An important avenue for future research is trust calibration, which ensures that users’ trust levels align with the model’s actual performance. Over-reliance on faulty outputs can lead to critical errors, while under-trusting a model may result in missed opportunity and efficiency gains. Future research could investigate feedback mechanisms that help users understand why a model generated certain outputs or offer insights into the decision-making process. Transparency tools, such as model explanations and uncertainty indicators, would allow users to better align their trust with the model’s reliability. This would be especially valuable in high-stakes SE tasks, where the consequences of errors can be significant.
@@
Summary of $RQ_{3}$: Trust Metrics We found that trust in LLM4SE is a multidimensional construct influenced by factors such as accuracy, usability, comprehensibility, and ethical compliance. A key finding from the user study revealed that 96% of the participants preferred multi-item trust metrics over single-item ones, emphasizing the need to capture trust comprehensively. User trust in LLMs for SE is influenced by both technical and human-centric factors. Accuracy and usability emerge as primary trust metrics, with accuracy being the most frequently cited criterion. Developers also highlighted the importance of comprehensibility and ethical compliance, particularly for tasks like program repair and test case generation. Sociological factors, such as the reputation of the model’s developers and its user base, further influence trust perceptions. The findings underscore the necessity of developing multi-dimensional trust measurement frameworks that account for both technical performance and user-centric dimensions to better evaluate and foster trust in LLMs.
@@ 8. Threats to Validity
We conducted our literature review following guidelines by Kithenham et al.  (Kitchenham, 2012), and our Survey Study using purposive sampling  (Baltes and Ralph, 2021). However, limitations may exist in our search strategy, data analysis, and participant recruitment.
External Threats. One potential threat to the generalizability of our results is our search strategy. Our search string might have missed some studies. To mitigate this threat, we tried multiple search strategy: 1) ”Large Language Model” OR ”Software Engineering” OR ”Trust”, 2) ”Software Engineering” OR ”SE” OR ”LLM” OR ”Large Language Model” OR ”Trust”, 3) ”Large Language Model” AND ”Software Enginnering” AND ”Trust”, 4) ”LLM” OR ”SE” OR ”Trust” OR ”Distrust” OR ”Trustworthiness”, 5) ”LLM” OR ”Large Language Model” OR ”SE” OR ”Software Engineering” OR ”Trust” OR ”Distrust” OR ”Trustworthiness”, 6) (“trust” OR “distrust” OR “trustworthiness”) AND (“SE” OR “Software Engineering”) AND (“LLM” OR “Large Language Model” OR “LLMs” OR “Large Language Models” OR “Deep Learning” OR “Machine Learning”), 7) (“trust” OR “distrust” OR “trustworthiness”) AND (“SE” OR “Software Engineering” OR “Software Development” OR “Code Generation” OR “Program Repair” OR ”Automatic Program Repair” OR “Software Testing” OR “Test Generation”) AND (“LLM” OR “Large Language Model” OR “LLMs” OR “Large Language Models” OR “Deep Learning” OR “Machine Learning”). Search string 7 returned the most results. Although this increased the time required to filter out unrelated papers, it reduced selection bias due to the large pool of results. Another threat might arise from our inclusion and exclusion criteria. To remove biases, we predefined our criteria and conducted the methodology independently by two authors. Additionally, we also conducted complementary analysis of $70$ papers that were filtered out from our exclusion criteria but contained trust concepts from broader literature.
While our user study consists of $25$ practitioners, it is important to note that this reflects a highly specialized group of individuals with direct experience in LLMs for SE tasks. Given the niche nature of LLM applications in SE, identifying qualified participants is inherently challenging. Similar SE studies have successfully used small, targeted samples (Krishna et al., 2024; coteallard2019deeplearningelectromyographichand, ). Additionally, filtering questions ensured participants were qualified, and only qualifying responses were used for analysis. We also published our responses in an online appendix (AnonymousRepoTrustRepo, 2024). While we acknowledge that larger-scale validation would be beneficial, this study provides an essential first step in mapping the trust landscape of LLMs in SE, laying the groundwork for future research.
@@ 9. Conclusion
The motivation for this study was centered on a fundamental question: Why is trust research essential in LLM4SE? To answer this, we conducted a comprehensive literature review, gathered practitioners’ feedback through a survey study, and analyzed broader trust literature. This approach allowed us to identify key gaps in how trust is defined, what factors shape it, and how it is measured. Our work provides a structured understanding of these concepts, mapping the current research landscape and highlighting their critical role in shaping LLM adoption in SE workflows.
Our findings reveal that trust in LLMs is neither absolute nor universal; it is highly task dependent and must be carefully calibrated. Developers who overtrust LLM-generated code risk introducing security vulnerabilities, code smells, and data leaks, while excessive distrust limits the potential benefits of automation and efficiency. We demonstrate that trust is influenced by multiple factors, including accuracy, robustness, interpretability, and ethical considerations. However, trust alone is not enough—trustworthiness, as an inherent property of an LLM, must be systematically evaluated through well-defined metrics rather than inferred from user perceptions alone.
By highlighting the need for structured trust assessment frameworks, our study lays the groundwork for future research on trust-aware LLM integration. As LLMs continue to shape modern software development, ensuring that trust is well-calibrated will be essential for maximizing their benefits while mitigating risks. Moving forward, researchers should focus on developing standardized trust metrics and improving transparency mechanisms to align LLM capabilities with developer expectations, ensuring their responsible and effective adoption in SE.

### v2

@@ 7.4. Discussion
Our study highlights the need for a comprehensive framework to measure trust in LLM4SE, as existing approaches do not account for its multifaceted nature. Although technical metrics such as accuracy, robustness, and accountability are crucial, our survey results indicate that trust also depends on usability, comprehensibility, and ethical considerations. Developers emphasized that trust is not solely determined by correctness, but also by how seamlessly LLMs integrate into their workflows. The need for minimal debugging and rework reinforces the importance of trust metrics that capture practical usability alongside model performance.
The survey findings also emphasize that the measurement of trust should be multidimensional. A strong preference (96% of participants) for multi-item trust metrics aligns with broader research indicating that single-item trust measures oversimplify complex trust dynamics. This is particularly critical in SE, where small errors can lead to security vulnerabilities and software failures. A multidimensional approach that incorporates accuracy, usability, ethical compliance, and consistency provides a more holistic assessment of trustworthiness.
Another major gap in current trust measurement frameworks is their inability to capture the temporal evolution of trust. Our findings indicate that trust in LLMs fluctuates over time, shaped by output quality and the extent to which developers feel they can rely on generated suggestions. This aligns with research in robotics and education, where trust is measured longitudinally through failure rates and behavioral consistency (Lee and See, 2004; Lewis et al., 2018). Future SE trust assessment frameworks should go beyond static evaluations and incorporate temporal metrics that track trust development over prolonged interactions with LLMs.
@@ 7.5. Opportunities for Future Work
The results of this study present several avenues for future research aimed at enhancing the metrics of trust in LLM4SE. One crucial avenue of future research based on our survey study is the development of a multi-item trust framework. We propose constructing a multi-item “LLM4SE Trustworthiness Index” that integrates objective performance metrics with user-centric evaluations. For example, based on our survey findings, key dimensions like Functional Correctness, Security & System Integrity, Understandability, and Usability would be prioritized. Operational examples for these metrics could include the following: Functional Correctness: Measured by automated test suite pass rates on LLM-generated code, or the percentage of generated code snippets that achieve desired functionality without modification. Security & System Integrity: Assessed by the rate of security vulnerabilities (e.g., using static analysis tools on generated code), or the frequency of nonmalicious yet harmful outputs. Understandability/Interpretability: Quantified through user ratings on code clarity, the time required for developers to comprehend and verify LLM-generated solutions for critical tasks or/and through interpretability techniques such as Code Rationale (Palacio et al., 2025), Shap (Lundberg and Lee, [n. d.]), and attention-based scores(Mohankumar et al., 2020). Usability/Workflow Integration: Evaluated using metrics such as reduction in manual debugging effort, time saved in task completion, or adherence to existing coding standards and project conventions. These individual metrics, drawn from both technical analysis and structured user feedback, would contribute to a composite trustworthiness score. Such an index could be task-specific, with different weightings for each dimension based on the task’s criticality (e.g., a higher weight for security in program repair vs. boilerplate generation). Furthermore, a robust framework would track these metrics longitudinally, allowing real-time calibration of user trust based on observed performance over time and across diverse contexts.
To effectively establish and validate such a comprehensive metric, a key future direction involves conducting causal analysis to determine the precise relative importance and interdependencies of each contributing factor. Understanding how different dimensions of trustworthiness (e.g., accuracy, interpretability, security) causally influence overall user trust will enable the development of more targeted and impactful trust-building interventions.
@@
Trust in LLMs evolves over time, influenced by repeated interactions and exposure to the model. Future research should explore longitudinal studies that track how trust builds or erodes as users gain experience with the model. These studies could examine changes in trust based on output quality, model performance over time, and user feedback. Understanding the temporal dynamics of trust would help in designing models that foster long-term user confidence, making them more effective and reliable for continuous use in SE tasks. Additionally, collecting extensive user-centric data from a larger, more diverse population will be necessary to ensure the generalizability and robustness of these measurements across various SE contexts and practitioner demographics. This scaled data collection will be crucial for validating proposed metrics and understanding how trust dynamics vary across different expertise levels, organizational cultures, and geographical locations.
@@
Context-specific trust metrics are another crucial area for future work. Trust may be prioritized differently in tasks such as code generation, program repair, or test case generation, where each task involves distinct expectations and challenges. For example, code generation can emphasize accuracy, ethical considerations, and contextual relevance, while program repair may require metrics that account for error minimization, maintainability, and adherence to coding standards. Tailored trust metrics will ensure that the measurement framework is aligned with the specific needs of each task and provides more meaningful insights into trust.
@@
An important avenue for future research is trust calibration, which ensures that users’ trust levels align with the model’s actual performance. Over-reliance on faulty outputs can lead to critical errors, while undertrusting a model may result in missed opportunities and efficiency gains. Future research could investigate feedback mechanisms that help users understand why a model generated certain outputs or offer insights into the decision-making process. Transparency tools, such as model explanations and uncertainty indicators, would allow users to better align their trust with the reliability of the model. This would be especially valuable in high-stakes SE tasks, where the consequences of errors can be significant.
Future work should therefore focus on developing sophisticated feedback mechanisms that empower developers to effectively interpret LLM outputs, moving beyond mere acceptance or rejection. This involves designing tools that provide actionable insights into the LLM’s reasoning and confidence levels, allowing practitioners to understand why a particular output was generated and how reliable it is for their specific task. Developing concrete recommendations for SE practitioners and tool builders regarding dynamic trust adjustment is crucial. For practitioners, this implies establishing heuristics for increasing scrutiny (e.g., for outputs in high-risk tasks or when confidence indicators are low) versus increasing reliance (e.g., for boilerplate code or outputs consistently meeting high-performance benchmarks). Tool builders should focus on integrating context-aware transparency tools and uncertainty indicators directly into SE environments, enabling dynamic trust alignment based on observed metrics.
@@
Summary of $RQ_{3}$: Trust Metrics We found that trust in LLM4SE is a multidimensional construct influenced by factors such as accuracy, usability, comprehensibility, and ethical compliance. A key finding from the user study revealed that 96% of the participants preferred multi-item trust metrics over single-item ones, emphasizing the need to capture trust comprehensively. User trust in LLMs for SE is influenced by both technical and human-centric factors. Accuracy and usability emerge as primary trust metrics, with accuracy being the most frequently cited criterion. Developers also highlighted the importance of comprehensibility and ethical compliance, particularly for tasks like program repair and test case generation. Sociological factors, such as the reputation of the model’s developers and its user base, further influence trust perceptions. The findings underscore the necessity of developing multidimensional trust measurement frameworks that take into account both technical performance and user-centric dimensions to better evaluate and foster trust in LLMs.
@@ 8. Threats to Validity
We conducted our literature review following the Kithenham et al. guidelines (Kitchenham, 2012) and our survey study using purposive sampling (Baltes and Ralph, 2021). However, limitations may exist in our search strategy, data analysis, and participant recruitment.
External Threats. One potential threat to the generalizability of our results is our search strategy. Our search string might have missed some studies. To mitigate this threat, we tried multiple search strategy: 1) ”Large Language Model” OR ”Software Engineering” OR ”Trust”, 2) ”Software Engineering” OR ”SE” OR ”LLM” OR ”Large Language Model” OR ”Trust”, 3) ”Large Language Model” AND ”Software Enginnering” AND ”Trust”, 4) ”LLM” OR ”SE” OR ”Trust” OR ”Distrust” OR ”Trustworthiness”, 5) ”LLM” OR ”Large Language Model” OR ”SE” OR ”Software Engineering” OR ”Trust” OR ”Distrust” OR ”Trustworthiness”, 6) (“trust” OR “distrust” OR “trustworthiness”) AND (“SE” OR “Software Engineering”) AND (“LLM” OR “Large Language Model” OR “LLMs” OR “Large Language Models” OR “Deep Learning” OR “Machine Learning”), 7) (“trust” OR “distrust” OR “trustworthiness”) AND (“SE” OR “Software Engineering” OR “Software Development” OR “Code Generation” OR “Program Repair” OR ”Automatic Program Repair” OR “Software Testing” OR “Test Generation”) AND (“LLM” OR “Large Language Model” OR “LLMs” OR “Large Language Models” OR “Deep Learning” OR “Machine Learning”). Search string 7 returned the most results. Although this increased the time required to filter out unrelated papers, it reduced selection bias due to the large pool of results. Another threat might arise from our inclusion and exclusion criteria. To remove biases, we predefined our criteria and independently conducted the methodology by two authors. Furthermore, we also conducted a complementary analysis of $70$ articles that were filtered from our exclusion criteria, but contained trust concepts from the broader literature.
Although our user study consists of $25$ practitioners, it is important to note that this reflects a highly specialized group of individuals with direct experience in LLMs for SE tasks. Given the niche nature of LLM applications in SE, identifying qualified participants is inherently challenging. Similar SE studies have successfully used small targeted samples (Krishna et al., 2024; Côté-Allard et al., 2019). Furthermore, filtering questions ensured that participants were qualified, and only qualifying responses were used for analysis. We also published our responses in an online appendix (AnonymousRepoTrustRepo, 2024). Although we acknowledge that larger-scale validation would be beneficial, this study provides an essential first step in mapping the trust landscape of LLMs in SE, laying the groundwork for future research. Future studies should aim to conduct larger surveys with more diverse samples, such as testers, developers, and managers, to enable more generalizable insights.
@@ 9. Conclusion
We center the motivation for this study around a fundamental question: Why is trust research essential in LLM4SE? To answer this, we conducted a comprehensive literature review, collected practitioners’ feedback through a survey study, and analyzed the broader trust literature. This approach allowed us to identify key gaps in the way trust is defined, what factors shape it, and how it is measured. Our work provides a structured understanding of these concepts, mapping the current research landscape, and highlighting their critical role in shaping LLM adoption in SE workflows.
Our findings reveal that trust in LLMs is neither absolute nor universal; it is highly task dependent and must be carefully calibrated. Developers who overtrust LLM-generated code risk introducing security vulnerabilities, code smells, and data leaks, while excessive distrust limits the potential benefits of automation and efficiency. We show that trust is influenced by multiple factors, including accuracy, robustness, interpretability, and ethical considerations. However, trust alone is not enough; trustworthiness, as an inherent property of an LLM, must be systematically evaluated through well-defined metrics rather than inferred from user perceptions alone.
By highlighting the need for structured trust assessment frameworks, our study lays the groundwork for future research on trust-aware LLM integration. As LLMs continue to shape modern software development, it will be essential to ensure trust is well calibrated to maximize their benefits while mitigating risks. Moving forward, researchers should focus on developing standardized trust metrics and improving transparency mechanisms to align LLM capabilities with developer expectations, ensuring their responsible and effective adoption in SE.
Our research shows that trust in LLM4SE is a multidisciplinary concept that requires collaborative efforts across fields to be fully understood and effectively operationalized. Insights from HCI can support the design of more intuitive trust affordances and interactive workflows. AI ethics can guide frameworks for fairness, accountability, and transparency in model behavior. Cognitive science can offer a deeper understanding of how trust is formed, calibrated, and disrupted in developer–LLM interactions. Incorporating these perspectives will help establish more robust and human-aligned approaches to trust in LLM-supported software engineering.